Problem 2 Determine the z parameters of the circuits below. 212 w 12 *** 392 W 492 WW
Answer #1

Given circuit SUI Z21 vh a V lo 211= ADMISSIONS CONTACTO CASA Zzz= 1 V Z cq vi ur 11 2+ 2+ 3+ 4 Zeq 7 X 7 7+ 1 8 Zeq=0.875 rVi = 8x3 3 38115 Zzi 221 = 0.62 V₂ Zzz Iz I=0 Iz 1 zea 7 11=0 V و با V2 = I2 X 2 222 222 = 3 2 1 1 20 - The Am Abode of LearnPy = 12 v=tly V = 12/1 Vi 212= Iz 2,2 zi Zi1=0.8752 Z21 -0.62 Zzz = 2 R 212=